The Singer Division of Patio Enterprises currently earns $3.92 million and has divisional assets of $24.5 million. The division manager is considering the acquisition of a new asset that will add to profit. The investment has a cost of $3,471,000 and will have a yearly cash flow of $864,000. The asset will be depreciated using the straight-line method over a six-year life and is expected to have no salvage value. Divisional performance is measured using ROI with beginning-of-year net book values in the denominator. The company’s cost of capital is 16 percent. Ignore taxes.

Required:

a. What is the divisional ROI before acquisition of the new asset? (Enter your answer as a percentage rounded to 1 decimal place (i.e., 32.1).)

b. What is the divisional ROI in the first year after acquisition of the new asset? (Enter your answer as a percentage rounded to 1 decimal place (i.e., 32.1).)
Answer #1

a) ROI before acquistion = Income / assets = 3.92/24.5= 16%

b) Annual depreciation = $3,471,000/6 = $578,500

Earnings from the asset = $864,000 - 578,500 = $285,500

ROI after acquisition of asset = (3920,000+285,500)/(24,500,000+3,471,000) = 15.0%